# Learning State-Invariant Representations of Objects from Image Collections with State, Pose, and Viewpoint Changes

### v1

@@ 4.1.2 Dual Embedding Space Method:
Object Loss: The loss $L_{piobj}$ identifies confusing instances between pairs of objects, shown as confusers in Fig. 4 (top). It pulls the confuser embeddings closer to the multi-image embedding of the same object, reducing intra-object distances (shown by green arrows), facilitating learning invariant features across transformations. Simultaneously, it increases inter-object distances by separating confuser embeddings and multi-image embeddings of different objects (shown by red arrows), aiding in learning discriminative features to distinguish between visually similar objects. This control over intra-object compactness and inter-object separability is achieved through margins $\alpha$ and $\beta$ respectively.
@@
Category Losses: The loss $L_{picat}$ clusters embeddings of objects from the same category together in the category embedding space as shown by the green arrows in Fig. 4 (bottom). It ensures that the distances between the single-image and multi-image embeddings of an object, and between the multi-image embeddings of different object-identities within the same category do not exceed a margin $\theta$. Conversely, the loss $L_{cat}$ utilizes the large-margin softmax loss [36] to separate embeddings of objects from different categories by a margin $\gamma$ in the category embedding space, as shown by the red arrows in Fig. 4 (bottom).
The $L_{picat}$ and $L_{piobj}$ losses optimize the distances between the embeddings of a pair of objects. PiRO sampled these pairs of objects randomly from the same category and trained the dual-encoder to disentangle the category and object representations in the dual embedding spaces by learning common features between them in the category space and distinctive features between them in the object space simultaneously.
4.2 Extending PiRO using Curriculum Learning
We now extend the PiRO framework [11] as follows to deal with the challenges caused by state changes of objects that exhibit similar visual characteristics: First, while PiRO utilizes a single-headed self-attention layer in its dual encoder in order to aggregate visual features from different viewpoints, our new approach uses multiple attention layers. This adaptation enables us to effectively capture invariant visual features under complex transformations beyond mere changes in pose. Secondly, we introduce a modified training approach that uses the same losses as [11], detailed in Sec. 4.1.2, but with a new curriculum learning strategy based on a novel mechanism for mining and joint training involving the different loss functions. While PiRO randomly samples a pair of objects from the same category, now we additionally sample similar objects from the same and other categories. As the training progresses, we sample objects that are harder to distinguish (based on the inter-object distances in the embedding space) to effectively learn discriminative object-identity representations in order to distinguish between visually similar objects.
4.2.1 A New Curriculum Learning Training Strategy:
@@
As shown in Fig. 4, the training begins by randomly sampling objects from the same category in the first epoch. Subsequently, at the start of each epoch, the learned multi-image object embeddings are used to identify similar objects from the same and other categories to guide the sampling process. Our method progressively samples objects that are closer in the learned embedding space as the training progresses.
Our training scheduler switches between the following three different sampling strategies to construct multi-image batches from pairs of objects for training:
@@
Similar to PiRO [11], it allows sampling diverse objects from the same category that may be visually different (such as a pink handbag and a yellow lunch box), helping the encoder learn common features between them in the category space and discriminative features in the object space.
@@
While the previous sampling strategy will include similar objects from the same category, the additional samples generated by this step increases the proportion of the visually similar pairs of objects from the same category in order to give greater weightage to such pairs. This allows the system to learn fine-grained differences between the similar-looking objects from the same category such as a pink handbag and a pink backpack from the category ‘bags’. At the start of the epoch, for each object in the dataset, a list of top-k neighboring objects in the learned object embedding space is retrieved, and, for pair formation, another object is randomly sampled from this list for training.
For (S1) and (S2), the dual encoder is jointly trained using the losses mentioned in Section 4.1.2, as follows
@@
At the start of the epoch, the object embedding space is partitioned into $f(N_{e})$ Voronoi cells using k-means, where $f$ is a function strictly increasing with the current epoch number $N_{e}$. Thus, the object embedding space is progressively divided into finer partitions as the training progresses. For each object in the dataset, another object is randomly sampled from the same partition. As training progresses, the pair of objects will be more difficult to separate in the object embedding space as the objects are going to be closer in finer partitions. This strategy allows sampling neighboring objects from any category that are visually similar in the object space helping the encoder learn discriminative features to distinguish between similar objects from any categories (such as a pink handbag and a pink T-shirt).
@@
where, $\mathcal{P}_{part}$ is the set of all object pairs where each pair comprises neighboring objects from the same partition in the learned object embedding space at the beginning of the epoch. Note that we do not use the $L^{x,y}_{picat}$ as the object pairs $(x,y)$ may be from different categories.
4.2.2 Training Details:
For a fair comparison with prior methods [7, 11], VGG-16 [37] is used as the CNN backbone. The last FC layers are modified to generate 2048-D embeddings and are initialized with random weights. Two-layer single-head self-attention layers are used for each embedding space with a dropout of 0.25. Training involves jointly optimizing the category and object-based losses, as described in Eqs. 1, 2. Margins $\alpha=0.25$ and $\beta=1.00$ are set for the object space, while margins $\theta=0.25$ and $\gamma=4.00$ are set for the category space, which are the same as [11]. $V$ is set to 12 images per object, and the images are resized to 224$\times$224 and normalized. We use the Adam optimizer with a learning rate of $5e^{-5}$. We train for 150 epochs and use the step scheduler that reduces the learning rate by half after every 30 epochs. For curriculum learning, the object embedding space is divided into $2N_{e}$ partitions, where $N_{e}$ is the current training epoch.
@@ 5 Experimental Results
5.1 Evaluation on ObjectsWithStateChange
TABLE II: Comparison of classification and retrieval performance for the category and object-based tasks against several state-of-the-art pose-invariant methods on the new ObjectsWithStateChange dataset. SV and MV denote either a single image or multiple images used at the time of inference respectively. The results indicate that our method outperforms prior methods on all tasks.
@@ 5.1 Evaluation on ObjectsWithStateChange
PI-CNN Single 82.57 84.29 25.37 26.74 54.74 84.75 86.87 11.35 36.89 54.96
PI-Proxy Single 84.51 86.86 27.10 28.85 56.83 86.69 89.33 12.27 39.83 57.03
PI-TC Single 81.55 84.29 42.04 52.12 65.00 84.21 88.52 22.02 61.72 64.12
PiRO Dual 88.07 92.30 69.01 74.78 81.04 89.44 94.16 68.97 81.86 83.61
Ours Dual 89.20 93.66 76.31 85.35 86.13 90.59 95.00 78.00 89.46 88.26
We compare our method with prior pose-invariant methods (PI-CNN, PI-Proxy, and PI-TC) by Ho et al. [7] that learn pose-invariant embeddings in the single embedding space. We also compare against the current state-of-the-art, PiRO [11] that learns pose-invariant category and object embeddings in two separate spaces, resulting in improved performance on object-level tasks.
@@
Figure 5: Comparison of the UMAP visualization of the embeddings of the 331 objects in our OWSC dataset learned using different methods. The embeddings for each object are shown using a unique color and pattern. The figures in the top row show the embeddings learned by the pose-invariant methods [7] that learn a single embedding (SE) space while the figures in the bottom row show the embeddings for methods learning a dual embedding (DE) space. It can be seen that the SE methods cannot separate objects from the same category. In comparison, the dual embedding (DE) methods can better separate the different objects despite the transformations. When comparing PiRO (bottom-left) and our proposed method (bottom-right), we see that the embeddings for different objects are better separated using our method.
@@
We train all of the prior methods on our transformation split and test on the probe split. The results are presented in Table II and we can observe that our method outperforms all the prior methods. This is because our method can separate fine-grained objects under different transformations more effectively than prior methods as demonstrated using UMAP visualization of the embeddings in Fig. 5. For our method, we can observe that the embeddings of each object are better clustered and separated from other objects. This is primarily due to the curriculum learning strategy that not only separates objects from the same category but also separates visually similar objects from other categories. Also, in the next subsection, we observe that using multiple attention layers further improves performance as it helps capture invariant features under complex transformations and state changes. We provide detailed ablation of both the curriculum learning and the architectural changes in Section 5.2. We further observe that our curriculum learning approach improves performance on object-level tasks on two other publicly available multi-view datasets (ObjectPI [7] and ModelNet-40 [1]) in Table III.
@@
(a) Results on single and multi-image invariant recognition and retrieval tasks on our ObjectsWithStateChange (OWSC) dataset.
@@
OWSC Same category 87.1 90.9 68.7 76.4 80.8 88.6 93.2 68.8 82.9 83.4
@@
Model Net40 Same category 85.2 88.9 93.7 96.9 91.2 79.7 86.1 84.0 98.2 87.0
@@
(a) Results on single and multi-image invariant recognition and retrieval tasks on our ObjectsWithStateChange (OWSC) dataset.
(b) Results on single and multi-view pose-invariant recognition and retrieval tasks on two publicly available multi-view datasets.
@@ 5.2 Ablation Studies
(A) Sampling Training Examples: PiRO uses multi-view images of a pair of objects randomly sampled from the same category, whereas our method additionally samples images of a pair of neighboring objects from the same category and other categories. We compare the performance of the two sampling strategies on the new ObjectsWithStateChange dataset on transformation-invariant recognition and retrieval tasks in Table III. We additionally also compare performance on pose-invariant recognition and retrieval tasks on two publicly available multi-view datasets (ObjectPI [7] and ModelNet-40 [1]) in Table III. For a fair comparison, the same architecture with a single-layer and single-head attention layer is used in both cases.
We observe that our strategy improves recognition and retrieval performance on object-based tasks. Specifically, in Table III, it improves single-view object recognition by 7.9% and single-view object retrieval by 9.2% over sampling objects randomly from the same category [11]. Similarly, in Table III, we observe that our sampling strategy improves single-view object recognition accuracy by 3.5% on ObjectPI and 1.8% on ModelNet, and single-view object retrieval mAP by 5.3% on ObjectPI and 3.9% on ModelNet. For multi-view object recognition and retrieval tasks, we see a significant improvement on our OWSC dataset and marginal improvements on ObjectPI and ModelNet-40 datasets.
However, we observe slight performance degradation on some of the category-based tasks. We conjecture that our curriculum learning approach focuses more on separating visually similar objects to improve performance on object-level tasks than category-level tasks. While these samples are harder to separate in the object embedding space, they are much easier to pull together in the category space as they have very similar attributes. Therefore, the samples that are informative for learning discriminative features between visually similar objects are not informative for learning common attributes between objects in the category space, thereby degrading performance on category tasks.
@@
(B) Architecture: As compared to [11] that used a single attention layer with a single head, we experimented with using multiple attention layers and multi-head attention to capture invariant visual features under complex transformations beyond just pose changes in the case of multi-view datasets. The results in Table IV indicate that using a single attention layer and multiple heads improves multi-image recognition and retrieval performance marginally. Using two attention layers with a single head yields the best performance. However, increasing the number of heads does not lead to any significant improvements for category-level tasks but yields marginal improvements on object-level tasks.
[…]
@@
(D) Retrieval Results: The single-image object retrieval results on our dataset presented in Fig. 7, show that our method can accurately retrieve images of the same object under various transformations in (a) on the left side and also effectively distinguish the correct object from similar ones in (b) on the right side. Figs. 8, 9 in the supplemental show more single-image retrieval results of visually similar objects from each category in our dataset.

### v2

@@ 4.2 Losses
Object Loss: As shown in Fig. 6 (top), $L_{piobj}$ reduces intra-object distances by pulling the embeddings of mutually confusing instances (confusers) between pairs of objects closer to the multi-image embedding of the same object (shown in green arrows), within a margin $\alpha$. Simultaneously, it increases the inter-object distances by separating the confusers and multi-image embeddings of different objects (shown in red arrows) by a margin $\beta$.
@@
Two Category Losses: As shown in Fig. 6 (bottom), $L_{picat}$ clusters the embeddings of pairs of objects within the same category (shown in green arrows), ensuring intra-category distances remain within a margin $\theta$. Meanwhile, $L_{cat}$ enforces inter-category separation by pushing the embeddings from different categories by a margin $\gamma$ (shown in red arrows).
Figure 7: This figure highlights the challenges caused by state changes. Significant variations in the appearance of an object in different states result in high intra-class variance among samples of the same object. More importantly, due to changes in state and pose, instances of distinct objects from the same and different categories can appear similar, making them hard to distinguish, as highlighted by the black bounding boxes. The figure includes examples from our OWSC dataset, such as a back pillow in its original state and another foldable pillow in a folded state, a chopped tomato and apple appearing similar, and so on (left). Notably, two red t-shirts (from the same category) and a red towel (from a different category) are indistinguishable when crumpled (right).
4.3 Using Curriculum Learning as a Mining Strategy
While the overall goal of the learning described in this paper is the same as in [11], its implementation is made complex by the fact that the objects are now allowed to undergo changes in their state. That is, we still want to make sure that the embeddings for the same object class are tightly clustered while remaining well separated from those for the other classes. However, that goal is difficult to achieve with a direct application of the PiRO framework [11] in the presence of all the additional variability injected by the changes in object state. Fig. 7 shows some examples of the similarities across classes as created by the changes in state.
To address these challenges created by the changes in state, we introduce a new curriculum learning strategy to mine informative samples for training. While PiRO randomly samples a pair of objects from the same category, now we additionally sample similar objects from the same and other categories. This ensures confusing instances of similar objects (as shown in Fig. 7) are adequately separated in the embedding space, regardless of their category. Furthermore, as the training progresses, we sample object pairs that are harder to distinguish to effectively learn discriminative object-identity representations in order to distinguish between visually similar objects. For this, we use the inter-object distances in the currently learned embedding space to identify confusing instances of objects (both within and across different categories) and separate them in the embedding space to facilitate learning distinctive features that remain consistent under state changes and other transformations.
We demonstrate later that our curriculum strategy helps better cluster images of objects undergoing state changes and separate confusing instances in Fig. 8, thereby leading to performance improvements, as compared to prior methods in Table III(b). This strategy facilitates learning more discriminative embeddings not only for datasets with state changes but also existing multi-view object datasets, as we will demonstrate later in in Table III(b).
4.3.1 Informative Sample Mining and Training:
This section explains our approach to train the dual-encoder model (ref. Sec. 4.1) using the losses (ref. Sec. 4.2), as illustrated in Fig. 6. These losses optimize the intra-class and inter-class distances between pairs of objects and we provide a detailed explanation of how these object pairs are selected next.
@@ 4.3.1 Informative Sample Mining and Training:
Training begins by randomly sampling object pairs from the same category in the first epoch. For each object pair, $V$ images are randomly sampled per object to train the encoder. In subsequent epochs, the learned multi-image object embeddings are used to identify similar object pairs from both the same and different categories at the start of each epoch to guide the sampling process. As training progresses, our method gradually selects object pairs that are closer in the learned embedding space.
During training, our method switches between three different sampling strategies to select object pairs, as described below:
@@
Similar to PiRO [11], it allows sampling diverse objects from the same category that may be visually different such as a pink handbag and a yellow lunch box (both from the category ‘bags’), helping the encoder learn common features between them in the category space and discriminative features in the object space.
@@
While the previous sampling strategy will include similar objects from the same category, the additional samples generated by this step increases the proportion of the visually similar pairs of objects from the same category in order to give greater weightage to such pairs. This allows the system to learn fine-grained differences between the similar-looking objects from the same category such as a pink handbag and a pink backpack from the category ‘bags’. At the start of the epoch, for each object in the dataset, a list of top-k neighboring objects in the learned object embedding space is retrieved using ANN techniques [18], and, for pair formation, another object is randomly sampled from this list for training.
For (S1) and (S2), the dual encoder is jointly trained using the losses mentioned in Section 4.2, as follows
@@
The previous two sampling strategies primarily focused on selecting samples within the same category. However, objects from different categories – especially those undergoing state changes or transformations – can also exhibit similar visual characteristics. For example, a red top, towel, cloth bag, and blanket from different categories may appear indistinguishable when crumpled or folded (see Fig. 7). This strategy aims to sample such object pairs irrespective of their category. For this, it utilizes the inter-object distances in the learned embedding space to select neighboring objects across all categories, aiming to effectively separate confusing instances of similar objects caused by state changes and other transformations. Furthermore, as training advances, the strategy progressively samples harder-to-distinguish object pairs with lower inter-object distances, regardless of their category.
Specifically, at the start of the epoch, the object embedding space is partitioned into $f(N_{e})$ Voronoi cells using k-means clustering, where $f$ is a function that increases strictly with the current epoch number $N_{e}$. In our approach, we use a ramp function $f(N_{e})=cN_{e}$, with a constant slope $c$, a lower bound $n_{min}$, and an upper bound $n_{max}$, as detailed further in the next subsection. This leads to progressively dividing the object embedding space into finer partitions as the training advances. For each object in the dataset, another object is randomly sampled from within the same Voronoi cell. As the training progresses, partitions become finer, yielding sampled object pairs that are closer in the object embedding space and share more similar characteristics, making them harder to distinguish.
@@
where, $\mathcal{P}_{part}$ is the set of all object pairs where each pair $(x,y)$ comprises neighboring objects from the same partition in the learned object embedding space at the beginning of the epoch. Note that $L^{x,y}_{picat}$ which clusters the embeddings of a pair of objects from the same category together in the category space, is not used as the object pairs $(x,y)$ may be sampled from different categories.
4.3.2 Implementation Details:
For a fair comparison with prior methods [7, 11], VGG-16 is used as the CNN backbone. The last FC layers are modified to generate 2048-D embeddings and are initialized with random weights. Two-layer single-head self-attention layers are used for each embedding space with a dropout of 0.25. Training involves jointly optimizing the category and object-based losses, as described in Eqs. 1, 2. Margins $\alpha=0.25$ and $\beta=1.00$ are set for the object space, while margins $\theta=0.25$ and $\gamma=4.00$ are set for the category space, which are the same as [11]. $V$ is set to 12 images per object, and the images are resized to 224$\times$224 and normalized. We use the Adam optimizer with a learning rate of $5e^{-5}$. We train for 150 epochs and use the step scheduler that reduces the learning rate by half after every 30 epochs. For curriculum learning, at the beginning of each epoch an IVFFlat index is built using FAISS [18] using the currently learned multi-image object embeddings (with one aggregated embedding per object). For this, the quantizer used is FlatL2 and the object embedding space is divided into $f(N_{e})=max(min(2\times N_{e},n_{max}),n_{min})$ partitions, where $N_{e}$ is the current training epoch, $n_{min}=8$ and $n_{max}=100$.
@@ 5 Experimental Results
5.1 Evaluation using ObjectsWithStateChange
TABLE II: Comparison of classification and retrieval performance for the category and object-based tasks against several state-of-the-art pose-invariant methods using the new ObjectsWithStateChange dataset. SV and MV denote either a single image or multiple images used at the time of inference respectively. The results indicate that our method outperforms prior methods on all tasks.
@@ 5.1 Evaluation using ObjectsWithStateChange
PI-CNN [7] Single 82.57 84.29 25.37 26.74 54.74 84.75 86.87 11.35 36.89 54.96
PI-Proxy [7] Single 84.51 86.86 27.10 28.85 56.83 86.69 89.33 12.27 39.83 57.03
PI-TC [7] Single 81.55 84.29 42.04 52.12 65.00 84.21 88.52 22.02 61.72 64.12
PiRO [11] Dual 88.07 92.30 69.01 74.78 81.04 89.44 94.16 68.97 81.86 83.61
Ours Dual 89.39 93.36 77.57 87.77 87.02 90.70 94.56 79.22 91.17 88.91
We compare our method with various state-of-the-art pose-invariant methods using our OWSC dataset. The prior methods were originally designed to learn pose-invariant embeddings from multi-view images, but can be adapted to learn invariant embeddings from object image collections with diverse transformations from arbitrary views as they cluster multiple images of the same object together without needing explicit pose, viewpoint, or state information during training.
Single Embedding Space Methods: Ho et al. [7] proposed three pose-invariant methods PI-CNN, PI-Proxy, and PI-TC that focused primarily on learning category-specific embeddings, where the object-to-object variations are represented by variations in embedding vectors within the same embedding space. Specifically, these methods clustered single-view embeddings close to the multi-view embedding of the same object, that is clustered close to a learned proxy embedding of its category. However, the embeddings of different objects within the same category are not explicitly separated, yielding poor performance on object-based tasks.
@@
Figure 8: Comparison of the UMAP visualization of the embeddings of the 331 objects in our OWSC dataset learned using different methods. The embeddings for each object are shown using a unique color and pattern. The figures in the top row show the embeddings learned by the pose-invariant methods [7] that learn a single embedding (SE) space while the figures in the bottom row show the embeddings for methods learning a dual embedding (DE) space. It can be seen that the SE methods cannot separate objects from the same category. In comparison, the dual embedding (DE) methods can better separate the different objects despite the state changes and transformations. When comparing PiRO (bottom-left) and our proposed method (bottom-right), we see that the embeddings for different objects are better separated using our method.
Dual Embedding Space Methods: To address this, PiRO [11] designed a dual-encoder model that decouples category and object representations in separate embedding spaces using margin-based losses specifically designed for each embedding space. Our proposed framework also learns dual embeddings by extending PiRO [11] using a curriculum learning approach, as described in Sec. 4.
@@
We train these methods using our OWSC dataset and compare the results in Table II. We observe that our method outperforms all the prior methods. This is because our method can separate fine-grained objects under different state changes and other transformations more effectively than prior methods as demonstrated using UMAP visualization of the embeddings in Fig. 8. Despite the high intra-class variance and similarity across different object classes caused by state changes, we can observe that the embeddings of each object are better clustered and separated from other objects for our method. This is primarily due to the curriculum learning strategy that not only separates objects from the same category but also separates visually similar objects from other categories. Also, in the next subsection, we observe that using multiple attention layers further improves performance as it helps capture invariant features under complex transformations and state changes. We provide detailed ablation of both the curriculum learning and the architectural changes in Section 5.2. We further observe that our curriculum learning approach improves performance on object-level tasks for three other multi-view datasets (ObjectPI, ModelNet, and FG3D) in Table III(b).
@@
OWSC Same category 87.1 91.2 68.7 76.4 80.9 88.5 93.2 68.8 82.9 83.4
@@
(a) Results on single and multi-image state-invariant recognition and retrieval tasks on our ObjectsWithStateChange (OWSC) dataset.
@@
Model Net40 Same category 85.2 88.9 94.0 96.9 91.2 79.7 86.1 84.0 98.2 87.0
@@
FG3D Same category 79.0 81.8 83.1 91.5 83.9 68.1 74.4 73.0 95.5 77.8
Curriculum (Ours) 78.3 81.3 84.7 92.7 84.3 67.6 73.7 77.3 96.1 78.7
(b) Results on single and multi-view pose-invariant recognition and retrieval tasks on three publicly available multi-view datasets.
@@ 5.2 Ablation Studies
(A) Curriculum Learning: PiRO uses multi-view images of a pair of objects randomly sampled from the same category, whereas our new method additionally samples images of a pair of neighboring objects from the same category and other categories. The question now is whether the additional computational burden entailed by this sampling (ref. Appendix B) is worth the effort. This ablation is an answer to that question.
We compare the performance of the two sampling strategies using our ObjectsWithStateChange dataset on transformation-invariant recognition and retrieval tasks in Table III(b). We additionally also compare performance on pose-invariant recognition and retrieval tasks using three publicly available multi-view datasets (ObjectPI [7], ModelNet-40 [1], and FG3D [2]) in Table III(b). For a fair comparison, the same architecture in [11] with a single-layer and single-head attention layer is used in both cases.
We observe that our strategy improves recognition and retrieval performance on object-based tasks. Specifically, in Table III(b), it improves single-view object recognition by 7.9% and single-view object retrieval by 9.2% over sampling objects randomly from the same category [11]. Similarly, in Table III(b), we observe that our sampling strategy improves single-view object recognition accuracy by 3.5% on ObjectPI, 1.5% on ModelNet, and 1.6% on FG3D, and single-view object retrieval mAP by 5.3% on ObjectPI, 3.9% on ModelNet, and 4.3% on FG3D. For multi-view object recognition and retrieval tasks, we see a significant improvement on our OWSC dataset in Table III(b) and marginal improvements on ObjectPI, ModelNet-40, and FG3D datasets in Table III(b).
However, we observe slight performance degradation on some of the category-based tasks. We conjecture that our curriculum learning approach focuses more on separating visually similar objects to improve performance on object-level tasks than category-level tasks. While these samples are harder to separate in the object embedding space, they are much easier to pull together in the category embedding space as they have very similar attributes, and are hence not informative for learning category-specific features.
@@
(B) Architecture: As compared to [11] that used a single attention layer with a single head, we experimented with using multiple attention layers and multi-head attention to capture invariant visual features under complex transformations beyond just pose changes in the case of multi-view datasets. The results in Table IV indicate that using a single attention layer and multiple heads improves multi-image recognition and retrieval performance marginally. Using two attention layers with a single head yields the best performance. However, further increasing the number of heads does not lead to any significant improvements for category-level tasks but yields marginal improvements on object-level tasks.
(C) Optimization of Intra-Object and Inter-Object Distances During Training: In Fig. 10 (top), we plot the percentage of training samples that are informative ($\tau_{\text{\it info}}$ in red). A training sample is considered informative if the loss $L_{piobj}$ for that sample is non-zero. Being a margin-based loss, if the training samples of different objects are already well separated or if the training samples of the same object are already well clustered then the loss $L_{piobj}$ for that sample will be zero and the sample is uninformative. If we compare $\tau_{\text{\it info}}$ at any given epoch, the curriculum learning approach generates more informative samples than sampling objects from the same category, especially during the initial stages of training.
@@
(D) Retrieval Results: The single-image object retrieval results on our dataset presented in Fig. 10, show that our method can accurately retrieve images of the same object under various transformations in (a) on the left side and also effectively distinguish the correct object from similar ones in (b) on the right side. Figs. 11, 12 in the supplemental show more single-image retrieval results of visually similar objects from each category in OWSC.